# Cohen–Macaulay Complexes, Duality Groups, and the dualizing module of ${\rm{Out}}(F_{N})$

### v1

@@ 1.1.1. Duality of the mapping class group via Lefschetz duality
For ${\rm{GL}}(N,\mathbb{Z})$ and mapping class groups, the dualizing module can be described in terms of the homology of a particularly nice $G$-complex. For ${\rm{GL}}(N,\mathbb{Z})$ this $G$-complex is the rational Tits building [BS73], whereas for mapping class groups the dualizing module is isomorphic to the the homology of the curve complex $\mathcal{C}(\Sigma)$ [Har86]. Via Bieri–Eckmann duality, such concrete descriptions of dualizing modules have enabled a wealth of cohomological calculations [CFP12, CP17, CFP19, MPP21, PS22, BMP+22, BPS23].
@@ 1.3. New descriptions of dualizing modules
If the space $X$ has the same dimension as $G$, then there is a further possible description of the dualizing module in terms of global section group $\Gamma(X)$ of the local homology sheaf. Let ${\rm{Hom}}_{c}(\Gamma(X),\mathbb{Z})$ be the group of compactly determined dual homomorphisms $f\colon\Gamma(X)\to\mathbb{Z}$. (A homomorphism $f$ is compactly determined if there exists a compact set $K$ such that $f(s)=0$ for every section $s$ that vanishes on $K$.) We say that the local homology sheaf on $X$ is semistable if the inverse system $\{\Gamma(K)\}_{K}$ of sections of the local homology sheaf over finite subcomplexes satisfies the Mittag-Leffler condition (see Section 2.6).
@@ 1.5.1. An example of Atanasov
One does not need to leave the world of RAAGs to find an aspherical CM complex $Y$ such that $\pi_{1}(Y)$ is not a duality group. Atanasov [Ata14] studied the two dimensional situation, and found an aspherical local-homotopy-CM complex with fundamental group $\mathbb{Z}^{2}\ast\mathbb{Z}^{2}$ (see Figure 2).
@@ 1.5.2. The CM property and visible irreducibility
Atanasov’s complex $Y$ given in Figure 2 satisfies slightly stronger conditions than having no free faces: the complex is visibly irreducible in the sense of Louder–Wilton (cf. [Wil22, Section 4], or Section 4 below). This is a condition for a complex to have no ‘obvious’ simplifications, and plays a key role in Louder–Wilton’s proof that that one-relator groups with negative immersions are coherent [LW24]. In Theorem E, we show that visible irreduciblity can be described in terms of the local homology sheaf.
Theorem E.
Let $Y$ be 2-complex. The following are equivalent:
@@
Theorem E, combined with Atanasov’s example, shows that even with strong restrictions on the local homology sheaf, fundamental groups of aspherical CM complexes may not be duality groups.
@@ 1.7. The structure of this article
In Section 4 we prove Theorem E on the relationship between the local homology sheaf and the visible irreducibility condition of Louder–Wilton.
@@ 2.1. Combinatorial sheaves and cosheaves
Taking $C^{\bullet}_{\sigma}(X)={\rm{Hom}}(C_{\bullet}^{\sigma}(X),\mathbb{Z})$ gives dual cochain complexes which compute local cohomology. This gives a (graded) cosheaf $h^{*}$ on $X$, with corestriction maps $h^{\ast}(\tau>\sigma):h^{\ast}(\tau)\to h^{\ast}(\sigma)$. Typically for us $X$ will be a local homology CM complex and we will only be interested in $h^{n}$.
@@ 2.3. Sheaf cohomology and cosheaf homology
$C_{c}^{k}(X;\mathcal{F}):=\bigoplus_{\sigma\in X_{k}}\mathcal{F}(\sigma),$
@@ 3.1. Poset preliminaries
and $p_{\leq}$, $p_{\geq}$ are defined similarly. The join $P\ast Q$ of two posets has underlying set $P\cup Q$ with partial orders on $P$ and $Q$ given by restriction, with the additional relation $p<q$ for all $p\in P$, $q\in Q$. If $\sigma=p_{0}<p_{1}<\cdots<p_{k}$ is a $k$-simplex in $|P|$, then its link is given by the join
(4) $(p_{0})_{>}\ast(p_{0},p_{1})\ast\cdots\ast(p_{k-1},p_{k})\ast(p_{k})_{<}$
We use $h_{P}(p)$ to denote the height of an element $p\in P$, dropping the subscript where the overlying poset is understood. This is defined to be the dimension of $p_{\geq}$, or equivalently the length of the longest chain in $p_{>}$.
@@
A homotopy CM complex is also homology CM. Using (4), one can obtain the following characterisation of the homotopy CM property (we learned this from Section 2 of [Vog90], where it appears with slightly notation).
Proposition 3.2 ([Vog90]).
@@
  • $p_{>}$ is $(h(p)-1)$-spherical for all $p\in P$,
  • $(p,p^{\prime})$ is $(h(p)-h(p^{\prime})-2)$-spherical whenever $p<p^{\prime}$ in $P$, and
  • $p_{<}$ is $(n-h(p)-1)$-spherical for all $p\in P$.
@@
Let $f\colon P\to Q$ be a poset map. If $f^{-1}(q_{\geq})$ is contractible for all $q\in Q$, then $f$ is a homotopy equivalence. If $f^{-1}(q_{\leq})$ is contractible for all $q\in Q$, then $f$ is a homotopy equivalence.
@@ 3.2. Thickening the spine of Outer space
Let $i\geq k+1$. The graph $T_{k}/F_{N}$ is is obtained from $T_{i}/F_{N}$ by collapsing a subgraph. As $T_{i}\in\mathcal{X}_{N}$ and $T_{k}\not\in\mathcal{X}_{N}$, at least one component of this subgraph has nontrivial fundamental group. Therefore there exists $T_{k}<T_{i}^{\prime}<T_{i}$ such that $T_{i}^{\prime}$ is a core extension of $T_{k}$ – the collapsed edges of $T_{i}^{\prime}/F_{N}$ form a core graph (a graph with no valence one vertices). If each $T_{i}^{\prime}$ is chosen maximally, the map taking the chain $T_{k+1}<T_{k+2}<\cdots<T_{l}$ to $T_{k+1}^{\prime}<T_{k+2}^{\prime}<\cdots<T_{l}^{\prime}$ gives a monotone retraction from $\phi^{-1}(c)$ to the poset of core extensions of $T_{k}$. This is shown to be contractible in Proposition 7.2 of [BG20]. Therefore $\phi$ is a homotopy equivalence by the Fiber Lemma.
@@ 3.3. The thickenings $\mathcal{L}_{N}$ and $\mathcal{K}_{N}$ are homotopy CM
If $p\in\mathcal{L}_{N}$ and $p=T_{0}<T_{1}<\cdots<T_{k}$ then $h_{\mathcal{L}_{N}}(p)=k$. Equivalently, height is the same in both $\mathcal{L}_{N}$ and $\mathcal{FS}_{N}^{\prime}$. The same result holds for $\mathcal{K}_{N}$ and ${\mathcal{FS}^{r}_{N}}^{\prime}$.
@@
$(T_{k})<(T_{k-1}<T_{k})<\cdots<(T_{0}<T_{1}<\cdots<T_{k})$
in $p_{\geq}$ witnesses $h_{\mathcal{L}_{N}}(p)=k$. ∎
@@
Let $p\in\mathcal{L}_{N}$. If $p\in\partial\mathcal{L}_{N}$ then $p_{>}$ is contractible, otherwise $p_{>}$ is homeomorphic to a $(h(p)-1)$–dimensional sphere.
Let $p\in\mathcal{K}_{N}$. If $p\in\partial\mathcal{K}_{N}$ then $p_{>}$ is contractible, otherwise $p_{>}$ is homeomorphic to a $(h(p)-1)$–dimensional sphere.
@@
We give the proof for $\mathcal{L}_{N}$ and omit the proof for $\mathcal{K}_{N}$. Suppose $p\in\mathcal{L}_{N}$ and let $p=T_{0}<T_{1}<\cdots<T_{k}$. If $p\not\in\partial\mathcal{L}_{N}$, then $T_{0}$, and hence each blow-up $T_{i}$, belongs to $\mathcal{X}_{N}$. Hence every subchain of $p$ is an element of $\mathcal{X}_{N}$. It follows that $p_{>}$ is the poset of all proper subchains of $p$, so is isomorphic to the boundary of a $h(p)$–simplex.
If $p\in\partial\mathcal{L}_{N}$ then $T_{0}\in\partial\mathcal{FS}_{N}$. Let $T_{j}$ be the first element of the chain lying in $\mathcal{X}_{N}$. The subchains of $T_{j}<T_{j+1}<\cdots<T_{k}$ span a (full) simplex $\Delta$ in $p_{>}$. Furthermore, the monotone retraction $\chi:\mathcal{L}_{N}\to\mathcal{L}_{N}$ given in Proposition 3.6 restricts to a retraction $r\colon p_{>}\to\Delta$, which is a homotopy equivalence by Corollary 3.4. Hence $p_{>}$ is contractible. ∎
@@
Suppose $p,p^{\prime}\in\mathcal{L}_{N}$ with $p<p^{\prime}$. Then $(p,p^{\prime})$ is homeomorphic to a $(h(p^{\prime})-h(p)-2)$–dimensional sphere.
Suppose $p,p^{\prime}\in\mathcal{K}_{N}$ with $p<p^{\prime}$. Then $(p,p^{\prime})$ is homeomorphic to a $(h(p^{\prime})-h(p)-2)$–dimensional sphere.
@@
Again we give the proof only for $\mathcal{L}_{N}$. Suppose $p,p^{\prime}\in\mathcal{L}_{N}$. If $p^{\prime}=T_{0}<T_{1}<\ldots<T_{k}$ then $p=T_{i_{0}}<T_{i_{1}}<\ldots<T_{i_{l}}$, with $0\leq i_{j}\leq k$ for $1\leq j\leq l$. In this case, $h(p^{\prime})-h(p)=k-l$.
As there are $k-l$ elements of $p^{\prime}$ that are not in $p$, the chains between $p$ and $p^{\prime}$ are determined by proper subsets of these $k-l$ elements. As a poset this is isomorphic to the boundary of a $k-l-1$ simplex. Hence $(p,p^{\prime})$ is homeomorphic to a sphere of dimension $k-l-2=h(p^{\prime})-h(p)-2$. ∎
@@
Suppose $p\in\mathcal{L}_{N}$. Then $p_{>}$ is homeomorphic to a wedge of spheres of dimension $3N-5-h(p)$.
Suppose $p\in\mathcal{K}_{N}$. Then $p_{>}$ is homeomorphic to a wedge of spheres of dimension $3N-5-h(p)$.
@@
As before, we assume $p\in\mathcal{L}_{N}$ and will make it clear where the proof differs for $p\in\mathcal{K}_{N}$. Suppose $h(p)=k$, so that $p=T_{0}<T_{1}<\cdots<T_{k}$ for some sequence of trees in $\mathcal{FS}_{N}$. As $p_{>}$ consists of superchains of $p$, it is isomorphic to the barycentric subdivision of
(5) $(T_{0})_{<}\ast(T_{0},T_{1})\ast(T_{2},T_{3})\ast\cdots\ast(T_{k})_{>},$
where these intervals are in $\mathcal{FS}_{N}$ (each factor in this join corresponds to a place where new elements can be added to the chain). Let $E_{i}$ denote the number edges in the graph $T_{i}/F_{N}$. As $(T_{0})_{<}$ consists of trees made by collapsing some, but not all, of the edge orbits in $T_{0}$, it is homeomorphic to the boundary of an $E_{0}-1$ simplex, and is an ($E_{0}-2$)-sphere. Similarly, $(T_{i-1},T_{i})$ is isomorphic to the poset of proper subsets of the edge orbits in $T_{i}$ that are collapsed when passing to $T_{i-1}$, so is an $(E_{i}-E_{i-1}-2)$-sphere.
Let $V_{k}$ be the number of vertices of $T_{k}/F_{N}$. The poset $(T_{k})_{>}$ is equal to the poset of blow-ups of $T_{k}/F_{N}$, which is homotopy equivalent to a wedge of spheres of dimension $32N-V_{k}-$ 3 (when $p\in\mathcal{L}_{N}$ this is shown in the proof of [Vog90, Corollary 2.7], when $p\in\mathcal{K}_{N}$ this is [Vog90, Corollary 3.2]). Putting together all of the above using Equation (5), and noting that there are $k+2$ terms in the join, we have shown that $p_{>}$ is homotopy equivalent to a wedge of spheres of dimension equal to
@@
We again omit the proof for $\mathcal{K}_{N}$. The homotopy CM condition for $\mathcal{L}_{N}$ follows from Proposition 3.2 and the work above. Suppose $\sigma=p_{0}<p_{1}<\cdots<p_{k}$ is a simplex in $\mathcal{L}_{N}$. The link of $\sigma$ decomposes as a join as described in Equation (4). In the work above, these join components are always nontrivial wedge products of spheres unless $p_{0}\in\partial\mathcal{L}_{N}$, in which case $(p_{0})_{>}$ is shown to be contractible in Proposition 3.11. Therefore the link of $\sigma$ is homotopy equivalent to a nontrivial wedge of spheres unless $p_{0}\in\partial\mathcal{L}_{N}$, when the link is contractible. As $\partial\mathcal{L}_{N}$ is upwards-closed in $\mathcal{L}_{N}$ the element $p_{0}\in\partial\mathcal{L}_{N}$ if and only if $p_{i}\in\partial\mathcal{L}_{N}$ for all $i$, which is if and only if $\sigma\in\partial\mathcal{L}_{N}$. ∎

### v2

@@ 1.1.1. Duality of the mapping class group via Lefschetz duality
For ${\rm{GL}}(N,\mathbb{Z})$ and mapping class groups, the dualizing module can be described in terms of the homology of a particularly nice $G$-complex. For ${\rm{GL}}(N,\mathbb{Z})$ this $G$-complex is the rational Tits building [BS73], whereas for mapping class groups the dualizing module is isomorphic to the the homology of the curve complex $\mathcal{C}(\Sigma)$ [Har86]. Via Bieri–Eckmann duality, such concrete descriptions of dualizing modules have enabled a wealth of cohomological calculations [CFP12, CP17, CFP19, MPP21, PS22, BMP+24, BPS23].
@@ 1.3. New descriptions of dualizing modules
If the space $X$ has the same dimension as $G$, then there is a further possible description of the dualizing module in terms of global section group $\Gamma(X)$ of the local homology sheaf. Let ${\rm{Hom}}_{c}(\Gamma(X),\mathbb{Z})$ be the group of compactly determined dual homomorphisms $f\colon\Gamma(X)\to\mathbb{Z}$. (A homomorphism $f$ is compactly determined if there exists a compact set $K$ such that $f(s)=0$ for every section $s$ that vanishes on $K$.) We say that the local homology sheaf on $X$ is semistable if the inverse system $\{\Gamma(K)\}_{K}$ of sections of the local homology sheaf over finite subcomplexes satisfies the Mittag-Leffler condition (see Section 2.6). Recall that while dualizing modules are always torsion-free, they are not necessarily free as abelian groups (see also Question 4).
@@ 1.4. Outer space and ${\rm{Out}}(F_{N})$
1.5. A bound for the top-dimensional cohomology of ${\rm{Out}}(F_{N})$
Bieri–Eckmann duality allows one to translate problems from high-dimensional cohomology to low-dimensional homology, at the cost of understanding the dualizing module. If $G$ is the mapping class group or ${\rm{GL}}(n,\mathbb{Z})$ and $d$ is the virtual cohomological dimension of $G$, then $H^{d}(G;\mathbb{Q})=0$. This is a result of Church–Farb–Putman [CFP12] in the case of the mapping class group, and Lee and Szczarba in the case of ${\rm{GL}}(n,\mathbb{Z})$ [LS76], and is obtained by a vanishing result for the coinvariants of the associated dualizing module. In the case of $G={\rm{GL}}(n,\mathbb{Z})$, the rational cohomology of $G$ also vanishes in codimensions one and two [CP17, BMP+24]. The analogous codimension-one vanishing result was shown not to hold in the mapping class group in work of Chan–Galatius–Payne [CGP21].
Infamously, one cannot prove any such vanishing result for ${\rm{Out}}(F_{N})$: the vcd of ${\rm{Out}}(F_{7})$ is 11 and Bartholdi [Bar16] showed that $H^{11}({\rm{Out}}(F_{7});\mathbb{Q})\cong\mathbb{Q}$. The proof is computer-based, and it is not clear if there is a geometric origin for this cohomology class. While we cannot solve this problem, the methods in this paper allow us to obtain a resolution of the dualizing module of ${\rm{Out}}(F_{N})$ via the chain complex for $H_{*}({\mathcal{X}_{N}^{r}},h^{2N-3}_{r})$. We describe this in Section 4, and use it to prove the following:
Theorem E.
Let $\rho\in{\mathcal{X}_{N}^{r}}$ be a rose in the spine of reduced outer space, and let ${\rm{Stab}}(\rho)<{\rm{Out}}(F_{N})$ be the stabilizer of this rose. Then
$\dim(H^{2N-3}({\rm{Out}}(F_{N});\mathbb{Q}))\leq\dim((h^{2N-3}(\rho)\otimes\mathbb{Q})_{{\rm{Stab}}(\rho)}),$
where $h^{2N-3}(\rho)$ is the local cohomology of the reduced spine at this rose.
A detailed investigation of $h^{2N-3}(\rho)$ as a ${\rm{Stab}}(\rho)$-representation is a topic for future research.
@@ 1.6.1. An example of Atanasov
One does not need to leave the world of RAAGs to find an aspherical CM complex $Y$ such that $\pi_{1}(Y)$ is not a duality group. Atanasov [Ata14] studied the two dimensional situation, and found an aspherical local-homotopy-CM complex (see Figure 2) with fundamental group $\mathbb{Z}^{2}\ast\mathbb{Z}^{2}$ which is not a duality group.
@@ 1.6.2. The CM property and visible irreducibility
Atanasov’s complex $Y$ given in Figure 2 satisfies slightly stronger conditions than having no free faces: the complex is visibly irreducible in the sense of Louder–Wilton (cf. [Wil24, Section 4], or Section 5 below). This is a condition for a complex to have no ‘obvious’ simplifications, and plays a key role in Louder–Wilton’s proof that that one-relator groups with negative immersions are coherent [LW24]. In Theorem F, we show that visible irreduciblity can be described in terms of the local homology sheaf.
Theorem F.
Let $Y$ be a 2-complex. The following are equivalent:
@@
Theorem F, combined with Atanasov’s example, shows that even with strong restrictions on the local homology sheaf, fundamental groups of aspherical CM complexes may not be duality groups.
@@ 1.8. The structure of this article
In Section 5 we prove Theorem F on the relationship between the local homology sheaf and the visible irreducibility condition of Louder–Wilton.
@@ 2.1. Combinatorial sheaves and cosheaves
Taking $C^{\bullet}_{\sigma}(X)={\rm{Hom}}(C_{\bullet}^{\sigma}(X),\mathbb{Z})$ gives dual cochain complexes which compute local cohomology. This gives a (graded) cosheaf $h^{*}$ on $X$, with corestriction maps $h^{\ast}(\tau>\sigma):h^{\ast}(\tau)\to h^{\ast}(\sigma)$. Typically for us $X$ will be a local homology CM complex and we will only be interested in the degree $n$ parts $h_{n}$ and $h^{n}$.
@@ 2.3. Sheaf cohomology and cosheaf homology
(1) $C_{c}^{k}(X;\mathcal{F}):=\bigoplus_{\sigma\in X_{k}}\mathcal{F}(\sigma),$
@@ 3.1. Poset preliminaries
and $p_{\leq}$, ${}_{\leq}p$ are defined similarly. The join $P\ast Q$ of two posets has underlying set $P\cup Q$ with partial orders on $P$ and $Q$ given by restriction, with the additional relation $p<q$ for all $p\in P$, $q\in Q$. If $\sigma=p_{0}<p_{1}<\cdots<p_{k}$ is a $k$-simplex in $|P|$, then its link is given by the join
(5) ${}_{<}(p_{0})\ast(p_{0},p_{1})\ast\cdots\ast(p_{k-1},p_{k})\ast(p_{k})_{<}$
We use $\textnormal{ht}_{P}(p)$ to denote the height of an element $p\in P$, dropping the subscript where the overlying poset is understood. This is defined to be the dimension of ${}_{\leq}p$, or equivalently the length of the longest chain in ${}_{<}p$.
@@
A homotopy CM complex is also homology CM. Using (5), one can obtain the following characterisation of the homotopy CM property (we learned this from Section 2 of [Vog90], where it appears with slightly different notation).
Proposition 3.2 ([Qui78]).
@@
  • ${}_{<}p$ is $(\textnormal{ht}(p)-1)$-spherical for all $p\in P$,
  • $(p,p^{\prime})$ is $(\textnormal{ht}(p)-\textnormal{ht}(p^{\prime})-2)$-spherical whenever $p<p^{\prime}$ in $P$, and
  • $p_{<}$ is $(n-\textnormal{ht}(p)-1)$-spherical for all $p\in P$.
@@
Let $f\colon P\to Q$ be a poset map. If $f^{-1}(_{\leq}q)$ is contractible for all $q\in Q$, then $f$ is a homotopy equivalence. If $f^{-1}(q_{\leq})$ is contractible for all $q\in Q$, then $f$ is a homotopy equivalence.
@@ 3.2. Thickening the spine of Outer space
Let $i\geq k+1$. The graph $T_{k}/F_{N}$ is obtained from $T_{i}/F_{N}$ by collapsing a subgraph. As $T_{i}\in\mathcal{X}_{N}$ and $T_{k}\not\in\mathcal{X}_{N}$, at least one component of this subgraph has nontrivial fundamental group. Therefore there exists $T_{k}<T_{i}^{\prime}<T_{i}$ such that $T_{i}^{\prime}$ is a core extension of $T_{k}$ – the collapsed edges of $T_{i}^{\prime}/F_{N}$ form a core graph (a graph with no valence one vertices). If each $T_{i}^{\prime}$ is chosen maximally, the map taking the chain $T_{k+1}<T_{k+2}<\cdots<T_{l}$ to $T_{k+1}^{\prime}<T_{k+2}^{\prime}<\cdots<T_{l}^{\prime}$ gives a monotone retraction from $\phi^{-1}(c)$ to the poset of core extensions of $T_{k}$. This is shown to be contractible in Proposition 7.2 of [BG20]. Therefore $\phi$ is a homotopy equivalence by the Fiber Lemma.
@@ 3.3. The thickenings $\mathcal{L}_{N}$ and $\mathcal{K}_{N}$ are homotopy CM
If $p\in\mathcal{L}_{N}$ and $p=T_{0}<T_{1}<\cdots<T_{k}$ then $\textnormal{ht}_{\mathcal{L}_{N}}(p)=k$. Equivalently, height is the same in both $\mathcal{L}_{N}$ and $\mathcal{FS}_{N}^{\prime}$. The same result holds for $\mathcal{K}_{N}$ and ${\mathcal{FS}^{r}_{N}}^{\prime}$.
@@
$(T_{1})<(T_{1}<T_{2})<\cdots<(T_{1}<\cdots<T_{k})$
in ${}_{<}p$ witnesses $\textnormal{ht}_{\mathcal{L}_{N}}(p)=k$. ∎
@@
Let $p\in\mathcal{L}_{N}$. If $p\in\partial\mathcal{L}_{N}$ then ${}_{<}p$ is contractible, otherwise ${}_{<}p$ is homeomorphic to a $(\textnormal{ht}(p)-1)$–dimensional sphere.
Let $p\in\mathcal{K}_{N}$. If $p\in\partial\mathcal{K}_{N}$ then ${}_{<}p$ is contractible, otherwise ${}_{<}p$ is homeomorphic to a $(\textnormal{ht}(p)-1)$–dimensional sphere.
@@
We give the proof for $\mathcal{L}_{N}$ and omit the proof for $\mathcal{K}_{N}$. Suppose $p\in\mathcal{L}_{N}$ and let $p=T_{0}<T_{1}<\cdots<T_{k}$. If $p\not\in\partial\mathcal{L}_{N}$, then $T_{0}$, and hence each blow-up $T_{i}$, belongs to $\mathcal{X}_{N}$. Hence every subchain of $p$ is an element of $\mathcal{X}_{N}$. It follows that ${}_{<}p$ is the poset of all proper subchains of $p$, so is isomorphic to the boundary of a $\textnormal{ht}(p)$–simplex.
If $p\in\partial\mathcal{L}_{N}$ then $T_{0}\in\partial\mathcal{FS}_{N}$. Let $T_{j}$ be the first element of the chain lying in $\mathcal{X}_{N}$. The subchains of $T_{j}<T_{j+1}<\cdots<T_{k}$ span a (full) simplex $\Delta$ in ${}_{<}p$. Furthermore, the monotone retraction $\chi:\mathcal{L}_{N}\to\mathcal{L}_{N}$ given in Proposition 3.6 restricts to a retraction $r\colon_{<}p\to\Delta$, which is a homotopy equivalence by Corollary 3.4. Hence ${}_{<}p$ is contractible. ∎
@@
Suppose $p,p^{\prime}\in\mathcal{L}_{N}$ with $p<p^{\prime}$. Then $(p,p^{\prime})$ is homeomorphic to a $(\textnormal{ht}(p^{\prime})-\textnormal{ht}(p)-2)$–dimensional sphere.
Suppose $p,p^{\prime}\in\mathcal{K}_{N}$ with $p<p^{\prime}$. Then $(p,p^{\prime})$ is homeomorphic to a $(\textnormal{ht}(p^{\prime})-\textnormal{ht}(p)-2)$–dimensional sphere.
@@
Again we give the proof only for $\mathcal{L}_{N}$. Suppose $p,p^{\prime}\in\mathcal{L}_{N}$. If $p^{\prime}=T_{0}<T_{1}<\ldots<T_{k}$ then $p=T_{i_{0}}<T_{i_{1}}<\ldots<T_{i_{l}}$, with $0\leq i_{j}\leq k$ for $1\leq j\leq l$. In this case, $\textnormal{ht}(p^{\prime})-\textnormal{ht}(p)=k-l$.
As there are $k-l$ elements of $p^{\prime}$ that are not in $p$, the chains between $p$ and $p^{\prime}$ are determined by proper subsets of these $k-l$ elements. As a poset this is isomorphic to the boundary of a $k-l-1$ simplex. Hence $(p,p^{\prime})$ is homeomorphic to a sphere of dimension $k-l-2=\textnormal{ht}(p^{\prime})-\textnormal{ht}(p)-2$. ∎
@@
Suppose $p\in\mathcal{L}_{N}$. Then $p_{<}$ is homeomorphic to a wedge of spheres of dimension $3N-5-\textnormal{ht}(p)$.
Suppose $p\in\mathcal{K}_{N}$. Then $p_{<}$ is homeomorphic to a wedge of spheres of dimension $3N-5-\textnormal{ht}(p)$.
@@
As before, we assume $p\in\mathcal{L}_{N}$ and will make it clear where the proof differs for $p\in\mathcal{K}_{N}$. Suppose $\textnormal{ht}(p)=k$, so that $p=T_{0}<T_{1}<\cdots<T_{k}$ for some sequence of trees in $\mathcal{FS}_{N}$. As $p_{<}$ consists of superchains of $p$, it is isomorphic to the barycentric subdivision of
(6) ${}_{<}(T_{0})\ast(T_{0},T_{1})\ast(T_{2},T_{3})\ast\cdots\ast(T_{k})_{<},$
where these intervals are in $\mathcal{FS}_{N}$ (each factor in this join corresponds to a place where new elements can be added to the chain). Let $E_{i}$ denote the number edges in the graph $T_{i}/F_{N}$. As ${}_{<}(T_{0})$ consists of trees made by collapsing some, but not all, of the edge orbits in $T_{0}$, it is homeomorphic to the boundary of an $E_{0}-1$ simplex, and is an ($E_{0}-2$)-sphere. Similarly, $(T_{i-1},T_{i})$ is isomorphic to the poset of proper subsets of the edge orbits in $T_{i}$ that are collapsed when passing to $T_{i-1}$, so is an $(E_{i}-E_{i-1}-2)$-sphere.
Let $V_{k}$ be the number of vertices of $T_{k}/F_{N}$. The poset $(T_{k})_{<}$ is equal to the poset of blow-ups of $T_{k}/F_{N}$, which is homotopy equivalent to a wedge of spheres of dimension $32N-V_{k}-$ 3 (when $p\in\mathcal{L}_{N}$ this is shown in the proof of [Vog90, Corollary 2.7], when $p\in\mathcal{K}_{N}$ this is [Vog90, Corollary 3.2]). Putting together all of the above using Equation (6), and noting that there are $k+2$ terms in the join, we have shown that ${}_{<}p$ is homotopy equivalent to a wedge of spheres of dimension equal to
@@
We again omit the proof for $\mathcal{K}_{N}$. The homotopy CM condition for $\mathcal{L}_{N}$ follows from Proposition 3.2 and the work above. Suppose $\sigma=p_{0}<p_{1}<\cdots<p_{k}$ is a simplex in $\mathcal{L}_{N}$. The link of $\sigma$ decomposes as a join as described in Equation (5). In the work above, these join components are always nontrivial wedge products of spheres unless $p_{0}\in\partial\mathcal{L}_{N}$, in which case ${}_{<}(p_{0})$ is shown to be contractible in Proposition 3.11. Therefore the link of $\sigma$ is homotopy equivalent to a nontrivial wedge of spheres unless $p_{0}\in\partial\mathcal{L}_{N}$, when the link is contractible. As $\partial\mathcal{L}_{N}$ is upwards-closed in $\mathcal{L}_{N}$ the element $p_{0}\in\partial\mathcal{L}_{N}$ if and only if $p_{i}\in\partial\mathcal{L}_{N}$ for all $i$, which is if and only if $\sigma\in\partial\mathcal{L}_{N}$. ∎